Raffles Bank pays compound interest of 4% per annum,calculated quarterly. If $2000 was deposited in the bank for a period of 46 years, calculate the interest earned in Raffles Bank.
thanks for help!

Answer by stanbon(75887)

Raffles Bank pays compound interest of 4% per annum,calculated quarterly. If $2000 was deposited in the bank for a period of 46 years, calculate the interest earned in Raffles Bank.
---------
A(t) = P(1+(r/n))^(nt)
-----
A(46) = 2000(1+(0.04/4))^(4*46)
A(46) = $12478.51
----
Interest earned was 12478.51-2000 = 10,478.51
==================================================
Cheers,
Stan H.